if a 6-sides die was rolled 3 times, what is the probability of getting
a) three 5's
b) no 5's
c) at least one 5
Answer #1

If one die rolled then possible outcomes is

s={1,2,3,4,5,6}

216 combination..when we rolled 3 die.

if die roll three times possible outcomes is

1) three fives

=first die+second die +3rd die

=1/6 *. 1/6 *. 1/6

=0.00463

Or 1/216

=0.00463

2)no fives

=5/6 * 5/6 * 5/6

=0.5787

3)43 combination of at least one 5s

=43/216

=0.20